# Dahyeon Kye, Changhyun Roh, Sukhun Ko, Chanho Eom, and Jihyong Oh https://github.com/CMLab-Korea/Awesome-Video-Frame-Interpolation

### v1

@@ 5.3.1 Image-level Metrics
Peak Signal-to-Noise Ratio (PSNR) quantifies reconstruction fidelity based on the mean squared error (MSE) between interpolated frame and GT frame. While higher PSNR reflects better numerical similarity, it often fails to align with human perception, especially for high-frequency or perceptually salient regions.
Structural Similarity Index (SSIM) [196] evaluates local structural integrity by comparing luminance, contrast, and texture patterns. SSIM values range in $[-1,1]$, with higher values indicating stronger structural alignment. Though more perceptually aligned than PSNR, SSIM may still overrate visually implausible outputs if global structure is preserved.
Interpolation Error (IE) [189] computes the root-mean-square error (RMSE) between interpolated frame and the GT frame. Despite being intuitive, IE shares limitations with PSNR in terms of perceptual relevance.
5.3.2 Perceptual Metrics
Perceptual metrics aim to assess the semantic plausibility, texture fidelity, and structural realism of interpolated frames, often aligning better with human visual preferences.
Natural Image Quality Evaluator (NIQE) [197] is a no-reference score derived from deviations to natural image statistics. Lower values reflect more natural, HQ frames.
Fréchet Inception Distance (FID) [198] measures the Fréchet distance between the feature distributions of generated frames and GT frames using a pre-trained Inception network [208]. Lower FID indicates better semantic alignment.
Learned Perceptual Image Patch Similarity (LPIPS) [199] measures perceptual similarity using deep features from pretrained networks. It is robust to minor misalignment and sensitive to semantic differences. Lower LPIPS signifies better perceptual quality.
FloLPIPS [200] extends LPIPS by applying motion-aware weighting based on optical flow. It emphasizes visual fidelity in regions undergoing large displacement.
STLPIPS [201] improves LPIPS by incorporating shift-tolerant feature matching, enhancing robustness to slight misalignments.
DISTS (Deep Image Structure and Texture Similarity) [202] separately evaluates texture and structure similarity using deep features. It balances local detail and global consistency.
5.3.3 Video-level Metrics
@@
VSFA [203] is a no-reference model trained on human labels. It estimates perceptual quality by aggregating deep features with a recurrent network. Lower scores suggest better perceived video quality.
@@
tOF [204] computes temporal optical flow consistency across frames. Lower tOF values indicate smoother motion continuity.
@@
Fréchet Video Distance (FVD) [205] measures the Fréchet distance between distributions of deep features extracted from real and generated videos using a pre-trained Inflated 3D ConvNet (I3D) [209]. Lower FVD values reflect stronger temporal and perceptual realism.
@@
Fréchet Video Motion Distance (FVMD) [206] improves upon FVD by disentangling motion and appearance, focusing more explicitly on dynamic consistency.
@@
VBench [207] is a multi-dimensional benchmark that scores video models across motion fidelity, coherence, and realism. It enables large-scale reference-free evaluation using semantic video representations.
6 Applications
6.1 Event-based VFI
Event-based Video Frame Interpolation (EVFI) [210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 14] aims to improve interpolation accuracy by leveraging the unique advantages of event cameras. Unlike conventional frame-based cameras that capture full images at fixed intervals, event cameras [224], which are bio-inspired vision sensors [225], asynchronously record per-pixel brightness changes, referred to as “events”, triggered when a contrast threshold is exceeded. These sensors offer key benefits such as ultra-high temporal resolution, high dynamic range, and low latency, making them ideal for scenarios involving rapid motion or challenging lighting. Consequently, event cameras have gained traction in VFI research, especially where traditional RGB frames suffer from motion blur or low temporal fidelity [212, 216, 218].
One of the early models, TimeLens [212], estimates optical flow directly from event streams and synthesizes intermediate frames accordingly. Later models such as TimeReplayer [216] and EGVD [223] improve performance by jointly estimating motion and appearance. TimeLens-XL [221] enhances any-time interpolation capability by optimizing flow and frame synthesis iteratively. Despite these advances, EVFI models remain sensitive to synthesis errors, as inaccuracies can accumulate over time, leading to temporal artifacts.
Despite their strengths, EVFI models face practical challenges. Capturing real event streams requires specialized neuromorphic sensors, which are often expensive and less accessible than conventional cameras. Moreover, collecting large-scale event datasets with dense GT labels is especially challenging due to the asynchronous nature of event recordings. As a result, several studies [226, 227, 228, 229] exploit event simulation from standard camera, simulating the event stream from continuous images or video sequences. Kaiser et al. [226] simulates positive or negative events by thresholding the intensity change between consecutive frames. Pix2NVS [227] estimates per-pixel luminance from video to synthesize event-like representations, aligned to frame intervals.
6.2 Cartoon VFI
Producing traditional 2D animation is labor-intensive [230], requiring artists to manually draw multiple in-between frames. VFI offers a means of automating this process by generating plausible intermediate frames, thereby reducing production time and cost [230, 231].
@@ 6.2 Cartoon VFI
However, cartoon videos exhibit distinct characteristics compared to real-domain videos: they feature exaggerated motion, minimal texture, flat color regions, and sharp contours, which pose challenges to correspondence-based methods. To address this, domain-specific models have been proposed [180, 232, 233, 109, 234, 235]. Notably, ToonCrafter [109] adopts a generative framework rather than relying on explicit motion estimation. Recent efforts aim to build models that generalize across both cartoon and real domains by leveraging diverse training data or domain adaptation techniques [114, 161, 116].
@@
A major bottleneck in cartoon VFI research is the absence of standardized, HQ datasets. While ATD-12K [180] provides a useful benchmark, its triplet-only format restricts its utility in ATFI settings. As a result, future progress will depend on the release of open, multi-frame cartoon datasets that enable fair and reproducible evaluation.
6.3 Medical Image VFI
@@
VFI is also increasingly applied in medical imaging to reconstruct temporally dense 4D sequences from sparsely acquired volumetric scans [236, 237, 238]. Modalities like CT and MRI face acquisition constraints due to radiation exposure and long scanning times [237], leading to coarse temporal sampling. VFI offers a means to generate intermediate volumes that enhance temporal resolution without incurring additional scan overhead. Medical VFI models must account for subtle anatomical motions and preserve fine structural detail critical for clinical interpretation. CPT-Interp [238] uses continuous motion field modeling, while DU4D [237] proposes an unsupervised interpolation framework that does not rely on GT annotations. These approaches enhance applicability in settings where labeled 4D medical datasets are scarce. Nonetheless, challenges remain. Ensuring clinical validity, minimizing hallucinated content, and establishing domain-specific evaluation metrics are ongoing concerns. Future research will likely explore physiology-aware modeling, uncertainty quantification, and benchmark design specific to 4D medical imaging tasks.
6.4 Joint Task
Recent studies have explored jointly performing VFI with other LLV tasks such as super-resolution (SR) [239, 240, 241, 8, 242, 243] and deblurring [244, 245, 246, 181, 247, 248]. Such joint formulations exploit the inherent correlation between spatial and temporal cues in video sequences [240]. For instance, space-time video super-resolution (STVSR) jointly upsamples resolution and frame rate by leveraging spatial details to enhance motion estimation and vice versa [240]. Shared representations enable efficient feature reuse, reduce redundancy, and facilitate joint optimization. Models such as FISR [240] and MOTIF [243] exemplify this integrated approach. Joint deblurring and interpolation addresses scenarios involving both motion blur and low frame rates. Instead of applying deblurring followed by VFI in a cascade, end-to-end models [139, 244, 245, 246, 181, 247, 248] simultaneously estimate clean and interpolated frames, resulting in improved temporal consistency and visual clarity. These multitask designs improve robustness and efficiency, particularly under real-world degradation, and suggest promising directions for unified LLV modeling.
7 Future Research Directions
7.1 Video Streaming Service
The widespread adoption of real-time video services, including video conferencing and adaptive streaming, presents a growing need for bandwidth-aware video delivery under constrained networks [249]. VFI offers a promising solution by enabling keyframe-only transmission while synthesizing intermediate frames on the client side, thus maintaining visual fluidity at lower bitrates. While early methods confirm its potential for rate reduction, practical deployment remains scarce due to model complexity, inference latency, and platform limitations. Future directions include the development of ultra-lightweight architectures that can operate on mobile or edge devices with limited compute resources. Moreover, adaptive interpolation strategies that jointly consider network bandwidth, scene motion complexity, and perceptual saliency are needed. Learning-based rate control, where the interpolation fidelity is dynamically modulated, could enable bitrate–quality trade-offs tuned in real-time. Joint optimization pipelines that integrate VFI models into codecs or reinforcement learning-based streaming agents may unlock robust low-latency video systems. In particular, methods that unify frame interpolation with residual-based encoding and decoding schemes could blur the boundary between generation and compression, laying the groundwork for next-generation streaming protocols.
7.2 All-in-One LLV Video Restoration
While all-in-one models have shown promising results in image restoration [250, 251, 252], equivalent progress in the video domain particularly in unified LLV frameworks, remains limited. Current LLV restoration pipelines remain fragmented, with VFI, denoising, deblurring, and SR often treated as separate tasks. This modularity, while convenient for controlled benchmarking, limits model robustness under real-world degradations that involve complex mixtures of temporal and spatial artifacts. A promising direction is the development of unified, all-in-one architectures that perform multiple LLV tasks jointly, where VFI is not treated as a standalone module but as an integral part of a broader restoration framework. The interpolated frames can offer temporally consistent guidance for denoising or deblurring, while super-resolved outputs can enhance motion estimation accuracy. Cross-task consistency losses or multi-task learning objectives can foster synergistic improvements. Moreover, transformer- or diffusion-based architectures with spatio-temporal attention mechanisms are naturally suited to this multi-task paradigm, as they can encode long-range dependencies and modulate task-specific pathways via conditioning.
7.3 3D and 4D Scene Understanding
VFI research remains largely grounded in 2D image space, often assuming planar motion and flat appearance fields. However, the increasing prevalence of 3D-aware applications in AR/VR, robotics, and multiview rendering calls for VFI methods that explicitly account for the underlying geometry of dynamic scenes. Recent works in 4D scene modeling using temporal neural fields [253], dynamic Gaussians [254], and neural point representations [255] suggest that temporally coherent synthesis is possible when motion is modeled in 3D space. Integrating VFI into such pipelines enables physically plausible interpolation that respects depth, occlusion, and parallax. Depth-conditioned flow, pose-aware synthesis, or geometry-aware latent spaces may serve as intermediate representations. Furthermore, VFI models can be extended to generate novel viewpoints, enabling geometry-consistent interpolation across spatial and temporal domains. Applications span from time-synchronized multiview interpolation to immersive scene reconstruction from sparse video inputs. Future work may explore co-training paradigms that fuse interpolation and view synthesis losses, jointly supervising geometry, appearance, and motion fields across space-time volumes.

### v3

@@ V-C1 Image-level Metrics
  • Peak Signal-to-Noise Ratio (PSNR $\uparrow$) quantifies reconstruction fidelity based on the mean squared error (MSE) between interpolated frame and GT frame. While higher PSNR reflects better numerical similarity, it often fails to align with human perception, especially for high-frequency or perceptually salient regions.
  • Structural Similarity Index (SSIM $\uparrow$) [234] evaluates local structural integrity by comparing luminance, contrast, and texture patterns. SSIM values range in [1,1][-1,1], with higher values indicating stronger structural alignment. Though more perceptually aligned than PSNR, SSIM may still overrate visually implausible outputs if global structure is preserved.
  • Interpolation Error (IE $\downarrow$) [5] computes the root-mean-square error (RMSE) between interpolated frame and the GT frame. Despite being intuitive, IE shares limitations with PSNR in terms of perceptual relevance.
V-C2 Perceptual Metrics
Perceptual metrics aim to assess the semantic plausibility, texture fidelity, and structural realism of interpolated frames, aligning with human visual preferences.
  • Natural Image Quality Evaluator (NIQE $\downarrow$) [160] is a no-reference score derived from deviations to natural image statistics.
  • Fréchet Inception Distance (FID $\downarrow$) [80] measures the Fréchet distance between the feature distributions of generated frames and GT frames using a pre-trained Inception network [213].
  • Learned Perceptual Image Patch Similarity (LPIPS $\downarrow$) [273] measures perceptual similarity using deep features from pretrained networks. It is robust to minor misalignment and sensitive to semantic differences.
  • FloLPIPS ($\downarrow$[45] extends LPIPS by applying motion-aware weighting based on optical flow. It emphasizes visual fidelity in regions undergoing large displacement.
  • STLPIPS ($\downarrow$[64] improves LPIPS by incorporating shift-tolerant feature matching, enhancing robustness to slight misalignments.
  • DISTS (Deep Image Structure and Texture Similarity $\downarrow$) [51] separately evaluates texture and structure similarity using deep features. It balances local detail and global consistency.
V-C3 Video-level Metrics
@@
  • VSFA ($\downarrow$[129] is a no-reference model trained on human labels. It estimates perceptual quality by aggregating deep features with a recurrent network.
@@
  • tOF ($\downarrow$[37] computes temporal optical flow consistency across frames.
@@
  • Fréchet Video Distance (FVD $\downarrow$) [223] measures the Fréchet distance between distributions of deep features extracted from real and generated videos using a pre-trained Inflated 3D ConvNet (I3D) [19]..
@@
  • Fréchet Video Motion Distance (FVMD $\downarrow$) [142] improves upon FVD by disentangling motion and appearance, focusing more explicitly on dynamic consistency.
@@
  • VBench $\downarrow$ [95] is a multi-dimensional benchmark that scores video models across motion fidelity, coherence, and realism.
TABLE IV: Quantitative comparison of DM-based methods on standard VFI benchmarks. We summarize the performance of representative methods on the DAVIS [184] benchmark. The results are reported in the order of PSNR $\uparrow$ / LPIPS $\downarrow$.
Methods DAVIS
Diffusion Model-based (§II-C)
MCVD [NeurIPS’22] [226] 18.946 / 0.247
LDMVFI [AAAI’24] [47] 19.98 / 0.276
PerVFI [CVPR’24] [240] 25.073 / 0.091
LBBDM [ACMMM’24] [152] 26.391 / 0.092
TRF [ECCV’24] [59] 14.132 / 0.484
VIDIM [CVPR’24] [102] 19.62 / 0.258
TLB-VFI [ICCV’25] [151] 26.27 / 0.086
GI [ICLR’25] [232] 14.850 / 0.246
ViBiDSampler [ICLR’25] [254] 14.811 / 0.448
FCVG [CVPR’25] [287] 16.162 / 0.247
TABLE V: Efficiency and complexity analysis of representative VFI models. We compare model size (Parameters $\downarrow$), computational cost (FLOPs $\downarrow$), and inference speed (Runtime $\downarrow$). The symbol ‘*’ denotes results measured by us using the official codes.
Method Parameters $\downarrow$ GFLOPs $\downarrow$ Runtime (Resolution) $\downarrow$ Hardware (GPU)
Kernel-based (§II-B1)
SepConv [ICCV’17] [170] 21.7 M 360 0.41 sec (1920 $\times$ 1080) NVIDIA RTX 3090
Flow-based (§II-B2)
DAIN [CVPR’19] [6] 24.0 M 5510 0.896 sec (640 $\times$ 480) NVIDIA RTX 3090
SoftSplat [CVPR’20] [168] 12.2 M 940 0.266 sec (1024 $\times$ 1024) NVIDIA RTX 2080Ti
ABME [ICCV’21] [178] 18.1 M 1300 1.16 sec (1920 $\times$ 1080) NVIDIA RTX 3090
XVFI [ICCV’21] [204] 5.6 M 370 0.36 sec (1920 $\times$ 1080) NVIDIA RTX 3090
RIFE [ECCV’22] [93] 9.8 M 200 0.035 sec (1024 $\times$ 1024) 2080 Ti
IFRNet [CVPR’22] [120] 5 M 210 0.10 sec (480 $\times$ 720) RTX A5000
AMT [CVPR’23] [133] 30.6 M 580 0.11 sec (480 $\times$ 720) NVIDIA RTX A5000
Transformer-based (§II-B6)
CAIN [AAAI’20] [36] 42.8 M 1290 0.14 sec (1920 $\times$ 1080) NVIDIA RTX 3090
VFIFormer [CVPR’22] [150] 24.1 M 47710 4.34 sec (480 $\times$ 720) NVIDIA RTX A5000
EMA-VFI [CVPR’23] [270] 65.6 M 910 0.70 sec (480 $\times$ 720) NVIDIA RTX A5000
Mamba-based (§II-B7)
VFIMamba [NeurIPS’24] [267] 66.1 M 940 0.23 sec (480 $\times$  850) NVIDIA V100
Diffusion Model-based (§II-C)
MCVD [NeurIPS’22] [226] 33.2 M* 349.6 G* 0.07 sec (256 $\times$ 256)* NVIDIA RTX 5090
LDMVFI [AAAI’24] [47] 461.6 M* 44.58 G* 1.34 sec (480 $\times$ 720)* NVIDIA RTX 5090
LBBDM [ACMMM’24] [152] 146.40 M* 1505 G* 3.21 sec (256 $\times$ 256)* NVIDIA RTX 5090
TRF [ECCV’24] [59] 2254 M* 106.6 G* 2.8 sec (1024 $\times$ 576)* NVIDIA RTX 5090
Framer [ICLR’25] [231] 1766 M* 365.9 G* 4.05 sec (512 $\times$ 320)* NVIDIA 5090
GI [ICLR’25] [232] 2254 M* 47.8 G* 2.69 sec (1024 $\times$ 576)* NVIDIA RTX 5090
ViBiDSampler [ICLR’25] [254] 2389 M* OOM 7.89 sec (1024 $\times$ 576)* NVIDIA A100
It is important to note that these evaluation metrics are not mutually independent, but instead reflect complementary and sometimes conflicting aspects of VFI quality. Image-level reconstruction metrics such as PSNR and SSIM [234] primarily favor pixel-wise fidelity and temporal averaging, often rewarding smooth predictions that minimize numerical errors. In contrast, perceptual metrics (e.g., LPIPS [273], FloLPIPS [45], FID [80]) emphasize semantic plausibility, texture sharpness, and high-frequency details, and may assign lower scores to outputs that are perceptually sharper but deviate locally from the GT. Video-level metrics further prioritize temporal stability and motion coherence across frames, penalizing flickering or inconsistent motion even when individual frames appear visually plausible.
As a result, optimizing for a single metric can inadvertently degrade performance along other dimensions. For example, models that aggressively smooth predictions to maximize PSNR may suffer from a loss of fine-grained motion details and over-smoothed edges that lead to visually washed-out textures, whereas generative or diffusion-based approaches that enhance perceptual realism may exhibit lower reconstruction scores due to their stochastic nature. This intrinsic trade-off highlights that VFI evaluation should be interpreted through the lens of metric intent and application context, rather than relying on any single indicator. A holistic assessment therefore benefits from jointly considering reconstruction fidelity, perceptual quality, and temporal consistency to better reflect real-world interpolation performance.
V-D Summary of Comparisons
To provide a comprehensive assessment of the VFI landscape, we present a systematic comparison of representative methods in terms of both quantitative result quality and computational efficiency. Table III provides a quantitative performance summary across standard benchmark datasets such as Vimeo-90K, Xiph-2K/4K, and SNU-FILM (Extreme). For each dataset, we report PSNR and LPIPS values. The results in Table III reveal distinct performance characteristics among different VFI methods. On the Vimeo-90K dataset, the transformer-based EMA-VFI [270] and Mamba-based VFIMamba [267] demonstrate strong reconstruction capabilities, achieving high PSNR values. Specifically, EMA-VFI attains competitive PSNR scores, reflecting the efficacy of attention mechanisms in capturing long-range dependencies for motion estimation. Similarly, VFIMamba and LC-Mamba [105] exhibit robust performance, indicating the potential of state space models in handling temporal dynamics efficiently. However, a notable trend emerges when comparing diffusion-based methods [226, 47, 152, 59, 231, 232, 254] with traditional deep learning approaches. While regression-based models (e.g., flow-based and transformer-based methods) generally yield higher PSNR values, they often struggle with perceptual quality, as evidenced by higher LPIPS scores. For instance, on Xiph-4K in Table III, the flow-based RIFE attains a PSNR of 33.760 dB with an LPIPS of 0.207, illustrating strong pixel-wise fidelity but only moderate perceptual quality when measured by learned perceptual distance.
In addition, we provide a focused analysis of diffusion-based VFI methods on the DAVIS [184] benchmark in Table IV. This table summarizes representative diffusion-based models [226, 47, 152, 59, 151, 287, 254, 240] and reports their performance on DAVIS in terms of PSNR and LPIPS. As shown in Table IV, recent approaches such as PerVFI and LBBDM achieve 25.073 dB / 0.091 and 26.391 dB / 0.092 (PSNR / LPIPS), respectively, that is, clearly lower LPIPS values at somewhat reduced PSNR compared with high-PSNR regression baselines. When interpreted together with the regression-based results, these trends highlight a characteristic trade-off of diffusion-based VFI: their stochastic generative formulation can yield sharper, more perceptually plausible in-between frames at the cost of slightly reduced pixel-wise fidelity. By reading Table IV alongside the broader benchmark in Table III, we make explicit that the relative strengths of diffusion-based methods become more apparent on perceptual indicators such as LPIPS, reinforcing the need to interpret quantitative results through the lens of metric intent rather than relying on PSNR alone.
Table V details the efficiency and complexity of representative VFI models, comparing model size (parameters), computational cost (GFLOPs), and inference runtime. The data reveals a significant trade-off between performance and computational resource requirements. Flow-based methods such as RIFE and IFRNet stand out for their efficiency, achieving real-time performance on standard resolutions with relatively low parameter counts and GFLOPs. These models are well-suited for applications where speed is critical. In contrast, transformer-based models such as VFIFormer and EMA-VFI, while offering high reconstruction accuracy, incur substantially higher computational costs, reflected in their larger parameter sizes and longer inference times. For example, EMA-VFI is built on heavy transformer backbones such as VFIFormer, whose complexity reaches approximately 47,710 GFLOPs, whereas Mamba-based models like VFIMamba achieve competitive reconstruction performance with only about 940 GFLOPs, as summarized in Table V. This contrast quantitatively illustrates that similar accuracy can be obtained at a fraction of the computational cost when adopting structured state-space backbones.
The disparity is even more pronounced for diffusion-based models. As shown in Table V, methods like LDMVFI, TRF, and Framer exhibit significantly higher GFLOPs and slower inference speeds compared to their regression-based counterparts. For example, TRF and GI require massive computational resources, with GFLOPs reaching into the thousands and runtime extending to several seconds per frame on high-end GPUs. This highlights the current bottleneck in adopting diffusion-based VFI for real-time applications, despite their perceptual advantages. Future research directions are therefore likely to focus on optimizing the sampling efficiency and architectural design of diffusion backbones to bridge this gap.
While no single algorithm achieves optimal results across all metrics and datasets, the choice of method depends heavily on the specific application requirements, that is, whether the priority lies in pixel-perfect reconstruction (favoring transformer/Mamba models), perceptual realism (favoring DM-based models), or real-time processing (favoring flow-based models). Moreover, these trade-offs are further modulated by input resolution and hardware capabilities, as models that operate in or near real time at moderate resolutions on high-end GPUs may become impractical for 4K content or resource-constrained devices, whereas lightweight architectures tend to maintain more stable throughput across such deployment scenarios. By concentrating this kind of quantitative evidence in Sec. V-D, we aim to provide readers with a rigorous, empirically grounded framework for interpreting the method characteristics and trade-offs discussed throughout the survey.
VI Applications
VI-A Event-based VFI
Event-based Video Frame Interpolation (EVFI) [235, 137, 222, 260, 275, 221, 79, 243, 118, 136, 146, 153, 24, 277, 215] leverages the unique properties of event cameras to enhance interpolation under fast motion and challenging lighting conditions. Unlike frame-based cameras, event cameras [135, 172] asynchronously record per-pixel brightness changes with ultra-high temporal resolution, high dynamic range, and low latency. These characteristics make them particularly effective when conventional RGB frames suffer from motion blur or insufficient temporal fidelity [222, 79, 118]. Early EVFI methods, such as TimeLens [222], synthesize intermediate frames by estimating motion directly from event streams. Subsequent works, including TimeReplayer [79] and EGVD [277], further improve performance through joint modeling of motion and appearance, while TimeLens-XL [153] extends EVFI toward any-time interpolation via iterative refinement. Despite these advances, EVFI remains sensitive to motion or event reconstruction errors, which can accumulate over time and cause temporal artifacts, highlighting the need for more robust event–frame fusion strategies.
Capturing real event streams requires specialized neuromorphic sensors, which are often more expensive and less accessible than conventional cameras. Moreover, collecting large-scale event datasets with dense GT labels is challenging due to the asynchronous nature of events. As a result, several studies [112, 12, 285, 278] simulate event streams from standard videos by modeling per-pixel intensity changes over time. From an application perspective, EVFI needs to balance latency and accuracy in streaming pipelines such as high-speed robotics or automotive perception. This motivates compact event encoders, on-device inference on neuromorphic hardware [193, 58], and hybrid designs that selectively fuse RGB frames and events based on motion intensity or lighting conditions.
VI-B Cartoon VFI
@@
Producing traditional 2D animation is labor-intensive [157], requiring artists to manually draw multiple in-between frames. VFI offers a means of automating this process by generating plausible intermediate frames, thereby reducing production time and cost [157, 22]. However, cartoon videos exhibit distinct characteristics compared to real-domain videos: they feature exaggerated motion, minimal texture, flat color regions, and sharp contours, which pose challenges to correspondence-based methods. To address this, domain-specific models have been proposed [208, 29, 131, 247, 256, 246]. Notably, ToonCrafter [247] adopts a generative framework rather than relying on explicit motion estimation. Recent efforts aim to build models that generalize across both cartoon and real domains by leveraging diverse training data or domain adaptation techniques [287, 231, 269].
@@
A major bottleneck in cartoon VFI is the absence of standardized, HQ datasets. While ATD-12K [208] provides a useful benchmark, its triplet-only format restricts its utility in ATFI settings. As a result, future progress will depend on the release of open, multi-frame cartoon datasets that enable fair and reproducible evaluation. In practical animation pipelines, deployment constraints include latency constraints for interactive editing tools, adaptation to varying line-art or shading styles, and model compression for integration into authoring software [208, 131]. These factors motivate style-robust architectures and lightweight backbones that can run in real-time or near real-time on commodity GPUs.
VI-C Medical Image VFI
@@
VFI is also increasingly applied in medical imaging to reconstruct temporally dense 4D sequences from sparsely acquired volumetric scans [72, 116, 130]. Modalities like CT and MRI face acquisition constraints due to radiation exposure and long scanning times [116], leading to coarse temporal sampling. VFI offers a means to generate intermediate volumes that enhance temporal resolution without incurring additional scan overhead. Medical VFI models must account for subtle anatomical motions and preserve fine structural detail critical for clinical interpretation. Methods like CPT-Interp [130] model continuous motion fields, and DU4D [116] performs unsupervised interpolation without GT labels, addressing the scarcity of annotated 4D datasets. Remaining challenges include ensuring clinical validity, preventing hallucinations, and developing domain-specific evaluation metrics. From an application standpoint, deployment requires strict control over hallucination risk, compatibility with existing reconstruction pipelines, and predictable latency for time-critical procedures. These requirements drive interest in uncertainty-aware VFI, physics- or deformation-constrained motion models, and memory-efficient architectures that can run within the hardware constraints of clinical scanners or PACS servers.
VI-D Joint VFI with LLV Task
Recent trends in video processing have moved toward unifying VFI with other LLV tasks, such as Super-Resolution (SR) [197, 117, 78, 244, 249, 30, 115] and deblurring [201, 200, 282, 174, 195, 257]. By exploiting the inherent correlation between spatial and temporal cues, these joint formulations achieve superior efficiency and performance compared to cascaded approaches. The most prominent joint task is Space-Time Video Super-Resolution (STVSR), which simultaneously increases spatial resolution and temporal frame rate [115, 244]. Unlike separate execution, STVSR models allow for shared feature representations, enabling efficient reuse of spatiotemporal information and joint optimization that reduces computational redundancy. Similarly, joint deblurring frameworks address scenarios where motion blur and low frame rates co-occur, such as in hand-held camera captures [201, 195]. Instead of applying deblurring followed by VFI sequentially, end-to-end models [276, 257] simultaneously estimate clean and interpolated frames. This holistic approach prevents error propagation from the deblurring stage, resulting in significantly improved temporal consistency and visual clarity. Extending these multitask capabilities, recent research has addressed high-noise environments, such as night-time surveillance, where video feeds suffer from extremely low Signal-to-Noise Ratio (SNR) [23]. In such regimes, conventional optical flow estimators fail to extract reliable correspondences, rendering standard “denoise-then-interpolate” cascades suboptimal. Consequently, joint VFI and restoration frameworks [195, 261] have emerged as a critical solution. By leveraging temporal redundancy to simultaneously hallucinate missing frames and suppress noise, these methods treat interpolation as a self-supervised restoration mechanism, ensuring robust motion modeling even under severe signal degradation. From a practical standpoint, joint LLV–VFI models are particularly attractive for video streaming and mobile applications where multiple enhancements must be executed under tight latency budgets. Current deployment trends focus on factorized architectures that reuse a common motion backbone across tasks, as well as model compression strategies to reduce memory footprint while maintaining high temporal fidelity.
VII Future Research Directions
VII-A Video Streaming Service
With the growth of real-time video services, bandwidth-efficient delivery has become critical. VFI can reduce transmission rates by sending only keyframes and generating intermediate frames on the client side, preserving smooth playback at lower bitrates. Key research directions include ultra-lightweight architectures for mobile and edge devices, and adaptive interpolation strategies that account for network bandwidth and scene motion. Integrating VFI into video codecs or streaming frameworks could enable robust, low-latency systems for next-generation video services. Beyond interpolation quality and efficiency, future deployment of VFI-enhanced video in streaming platforms will also require robustness against post-processing, screen-shooting, and malicious manipulation. Recent works [26, 28, 62, 132] have explored signal–noise separation for post-processed image forgery detection [26], flexible partial screen-shooting watermarking with provable robustness [28], wavelet-based screen-shooting watermarking and recovery [62], and grayscale-deviation-based modeling of screen-shooting distortion for robust watermarking [132]. Although these methods do not target VFI directly, they are complementary to VFI by helping secure interpolated content and model complex display–capture and compression channels in real-world video distribution.
VII-B All-in-One LLV Video Restoration
Although all-in-one architectures have achieved notable success in image restoration [128, 239, 1], their extension to unified low-level video (LLV) restoration remains underexplored. Existing pipelines typically decompose VFI, denoising, deblurring, and super-resolution into separate modules, which limits robustness under real-world degradations involving coupled spatial and temporal artifacts. A promising direction is the development of unified video restoration frameworks in which VFI is embedded as a core component rather than treated as an isolated task. In such settings, interpolated frames can provide temporally coherent priors for restoration, while improved spatial fidelity can in turn facilitate more reliable motion estimation. Multi-task learning objectives and cross-task consistency constraints may further promote synergy across tasks. Transformer- and diffusion-based architectures, with their strong spatio-temporal modeling capacity, are particularly well suited to this integrated paradigm.
VII-C 3D and 4D Scene Understanding
Most existing VFI methods operate purely in the 2D image space, implicitly assuming planar motion and appearance continuity. However, emerging applications in AR/VR, robotics, and multiview rendering demand interpolation frameworks that are aware of the underlying 3D scene structure. Recent advances in 4D scene representations, including temporal neural fields [180], dynamic Gaussian primitives [162], and neural point-based models [280], demonstrate that temporally coherent synthesis becomes more reliable when motion is modeled in 3D space. Incorporating VFI into such geometry-aware pipelines enables physically plausible interpolation that respects depth discontinuities, occlusions, and parallax effects. Promising directions include depth- or pose-conditioned interpolation, geometry-aware latent representations, and joint optimization schemes that unify frame interpolation and novel view synthesis. Such integration may allow VFI to evolve from a purely temporal task into a space-time consistent scene reconstruction component.
VII-D Physics-Informed VFI for Extreme Environments
Future VFI research must extend beyond standard scenarios to address extreme imaging conditions. A representative challenge is Underwater Imaging, which suffers from complex environmental degradations such as low visibility, light distortion, and color cast. These factors complicate the standard assumptions of brightness constancy and linear motion used in conventional VFI [216, 286, 2]. Drawing inspiration from the success of underwater image restoration, a promising direction is physics-informed VFI. By incorporating physical domain knowledge into deep learning pipelines, future models could better handle these unique distortions, improving temporal coherence for oceanographic and autonomous underwater applications.
VIII Conclusion
This survey reviewed the evolution of video frame interpolation (VFI) from classical motion-compensated methods to modern deep learning and generative approaches, and organized existing techniques through a unified taxonomy. While recent advances have significantly improved interpolation quality, fundamental challenges remain, including large motion, occlusion, lighting variation, and non-linear dynamics. Addressing these issues will require more accurate motion modeling, improved computational efficiency, and stronger generalization across diverse scenarios. We expect that continued integration of VFI with emerging video technologies will further broaden its applicability. This survey aims to serve as a concise reference and to motivate future research toward robust and efficient video synthesis.
Acknowledgments
This research was supported by the Chung-Ang University Research Scholarship Grants in 2024. This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2025-23524035). This research was supported by the MSIT(Ministry of Science and ICT), Korea, under the Graduate School of Virtual Convergence support program(IITP-2024-RS-2024-00418847) supervised by the IITP(Institute for Information & Communications Technology Planning & Evaluation)